2 vectors a and b have components in arbitrary units ax=3.2 ,ay=1.6, bx=0.50,by=4.5
(A)find angle between a and b 
(B)find the components of vector c that is perpendicular to  in xy plane  and has a megnitude of 5.0 units 

Grade:12th pass

1 Answers

Vikas TU
5 years ago
Dear Student,
Resultant a vector will be= sqroot(3.2^2 + 1.6^2)
                                = 3.5777
Resultant b vector will be = sqroot(0.5^2 + 4.5^2)
                                    = 4.527
Angle of a vector making woth horizontal, costhetha = 3.2/3.577 = 0.89
thetha = 26.5 degree.
 
Angle of b vector making with horizontal, costhetha = 0.5/4.527 = 0.1004
thetha = 84 degree.
Angle between a and b vector = 84 – 26.5 = 57.5 degree.